# DRL-Based Federated Self-Supervised Learning for Task Offloading and Resource Allocation in ISAC-Enabled Vehicle Edge Computing

### v1

@@ Conclusions
In this paper, we proposed an offloading task and resource allocation algorithm based on the SAC algorithm in Federated SSL. It employed this algorithm to allocate the transmission powers from vehicles to BS and RSU, CPU frequency for vehicle local training, assignment ratio of RSU computing resource. Based on the assignment ratio and the computing capability of RSU, this algorithm divided the local iterations task of vehicle into two parts: training locally and offloaded to the RSU. We also aimed to minimize the total energy consumption, including local training, RSU training, and transmission. Additionally, we set an offloading threshold. When the assignment ratio is below this threshold, the vehicle will not perform offloading operation. The conclusions in this paper are as follows:

### v2

@@ Conclusions
In this paper, we proposed an offloading task and resource allocation algorithm based on the SAC algorithm in federated SSL. It employed this algorithm to allocate the transmission powers from vehicles to BS and RSU, CPU frequency for vehicle local training, assignment ratio of RSU computing resource. Based on the assignment ratio and the computing capability of RSU, this algorithm divided the local iterations task of vehicle into two parts: training locally and offloaded to the RSU. We also aimed to minimize the total energy consumption, including local training, RSU training, and transmission. Additionally, we set an offloading threshold. When the assignment ratio is below this threshold, the vehicle will not perform offloading operation. The conclusions in this paper are as follows
@@
While the approach we suggested has shown promising results in smaller-scale IoV settings, it could encounter challenges related to scalability in larger or more intricate networks. In the future, we will further investigate this aspect to enhance the scalability of the scheme.